# Generalizable Spectral Embedding with an Application to UMAP

### v1

@@ Appendix G Technical Details
Table 4: Technical details of the real-world datasets used for GrEASE and NUMAP experiments.
Cifar10 Appliances KMNIST Parkinsons Wine Banknote
#samples 60,000 19735 70,000 5875 178 1372
#features 500 28 784 19 13 4
Table 5: Technical details in the GrEASE experiments for all datasets.
@@
To compute the ground truth SE on the train set and its corresponding eigenvalues, we constructed an affinity matrix $W$ from the train set (as detailed in Appendix C.2), with a number of neighbors detailed in Table 5. After constructing $W$, we computed the leading $k$ eigenvectors of its corresponding Unnormalized Laplacian $L=D-W$ via Python’s Numpy SVD or SciPy LOBPCG SVD (depending on the size). To get the generalization ground truth, we constructed an affinity matrix $W$ from the train and test sets combined, computed the leading $k$ eigenvectors of its corresponding Unnormalized Laplacian $L=D-W$, and extracted the representations corresponding to the test samples. We used a train-test split of 80-20 for all datasets.
@@
The architectures of GrEASE’s and SpectralNet’s networks in all of the experiments were as follows: size = 128; ReLU, size = 256; ReLU, size = 512; ReLU, size = $k+1$; orthonorm. NUMAP’s second NN and PUMAP’s NN architectures for all datasets was: size = 200; ReLU, size = 200; ReLU, size = 200; ReLU, size = 2; The SE dimensions for NUMAP were: Cifar10 - 20; Appliances - 10; Wine - 10; Banknote - 3.
The learning rate policy for GrEASE and SpectralNet is determined by monitoring the loss on a validation set (a random subset of the training set); once the validation loss did not improve for a specified number of epochs, we divided the learning rate by 10. Training stopped once the learning rate reached $10^{-7}$. In particular, we used the following approximation to determine the patience epochs, where $n$ is the number of samples and $m$ is the batch size: if $\frac{n}{m}\leq 25$, we chose the patience to be 10; otherwise, the patience decreases as $\max{(1,\frac{250m}{n})}$ (i.e., the number of iterations is the deciding feature).
To run UMAP, we used Python’s umap-learn implementation (UMAP’s formal implementation). We used the built-in initialization option ”spectral” (i.e., SE), and initialized contumely with PCA (implemented via Python’s sklearn.decomposition.PCA) and GrEASE. For Parametric UMAP we used the Pytorch implementaion (Liu, 2024). For all methods we used a default choice of 10 neighbors.

### v2

@@ Appendix G Technical Details
Table 11: Technical details of the real-world datasets used for Sep-SpectralNet and NUMAP experiments.
Cifar10 Appliances KMNIST Parkinsons Wine Banknote MNIST FashionMNIST
#samples 60,000 19735 70,000 5875 178 1372 60,000 60,000
#features 500 28 784 19 13 4 784 784
Table 12: Technical details in the Sep-SpectralNet experiments for all datasets.
@@
To compute the ground truth SE on the train set and its corresponding eigenvalues, we constructed an affinity matrix $W$ from the train set (as detailed in Appendix C.2), with a number of neighbors detailed in Table 12. After constructing $W$, we computed the leading $k$ eigenvectors of its corresponding Unnormalized Laplacian L=DWL=D-W via Python’s Numpy SVD or SciPy LOBPCG SVD (depending on the size). To get the generalization ground truth, we constructed an affinity matrix $W$ from the train and test sets combined, computed the leading $k$ eigenvectors of its corresponding Unnormalized Laplacian L=DWL=D-W, and extracted the representations corresponding to the test samples. We used a train-test split of 80-20 for all datasets.
@@
The architectures of Sep-SpectralNet’s and SpectralNet’s networks in all of the experiments were as follows: size = 256; ReLU, size = 256; ReLU, size = 512; ReLU, size = $k+1$; orthonorm. NUMAP’s second NN and PUMAP’s NN architectures for all datasets was: size = 200; ReLU, size = 200; ReLU, size = 200; ReLU, size = 2; The SE dimensions for NUMAP were: Cifar10 - 10; Appliances - 5; Wine - 10; Banknote - 3; Mnist - 10, FashionMnist - 10. For the datasets in Fig. 1, from top to bottom: Circles - 5, Cylinders - 11, Line - 2.
The learning rate policy for Sep-SpectralNet and SpectralNet is determined by monitoring the loss on a validation set (a random subset of the training set); once the validation loss did not improve for a specified number of epochs, we divided the learning rate by 10. Training stopped once the learning rate reached $10^{-7}$. In particular, we used the following approximation to determine the patience epochs, where $n$ is the number of samples and $m$ is the batch size: if $\frac{n}{m}\leq 25$, we chose the patience to be 10; otherwise, the patience decreases as $\max{(1,\frac{250m}{n})}$ (i.e., the number of iterations is the deciding feature).
To run UMAP, we used Python’s umap-learn implementation (UMAP’s formal implementation). We used the built-in initialization option "spectral" (i.e., SE), and initialized contumely with PCA (implemented via Python’s sklearn.decomposition.PCA) and Sep-SpectralNet. For Parametric UMAP we used the Pytorch implementaion (Liu, 2024). For all methods we used a default choice of 10 neighbors.
@@
For the computation of the p-values in Tab. 3 we used an independent t-test to compare the means of the 10 GS results, that were obtained from 10 runs using 10 different seeds.